# Rigorous Test for Quantum Integrability and Nonintegrability

### v2

@@ Abstract
Whether or not a quantum many-body system is integrable, which is characterized by the presence or absence of local conserved quantities, drastically impacts the dynamics of isolated systems, including thermalization. Nevertheless, a rigorous and comprehensive method for deciding integrability or nonintegrability has remained elusive. In this paper, we address this challenge by introducing rigorously provable tests for integrability and nonintegrability of quantum spin systems with finite-range interactions. Our approach establishes a new paradigm, moving beyond the conventional artisanal methods in the study of nonintegrability. Furthermore, it partially resolves the long-standing conjecture that integrability is governed by the presence or absence of local conserved quantities with small supports. The proposed framework guarantees that the nonintegrability of one-dimensional spin systems with translational symmetry can be confirmed algorithmically, regardless of system size.
@@ I Introduction
A central problem in quantum many-body physics is to understand the equilibrium states and dynamics of strongly interacting quantum many-body systems. Rigorous approaches sometimes provide deep insights into such puzzles. One direction is to establish physically meaningful results for broad classes of systems, such as the Hohenberg–Mermin–Wagner theorem [1, 2] and the Lieb–Schultz–Mattis theorem [3]. Another approach is the use of exactly solvable systems. The Bethe ansatz method, which originates from Bethe’s exact solution of the Heisenberg chain [4], has made it possible to obtain quantities that cannot analytically be calculated for general systems, such as energy eigenvalues, eigenstates, and correlation functions. Furthermore, the algebraic Bethe ansatz method, or the quantum inverse scattering method [5], has revealed that behind such solvability lies integrability, i.e., the existence of an extensive number of local conserved quantities.
@@
However, when considering the dynamics of isolated quantum many-body systems, integrable systems sometimes exhibit anomalous behavior. A typical example appears in thermalization of isolated quantum many-body systems [6]. Recent experiments have shown thermalization, i.e., the relaxation of local observables to thermal equilibrium in well-isolated quantum systems [7]. The eigenstate thermalization hypothesis (ETH) [8, 9, 10], which states that energy eigenstates are locally indistinguishable from thermal equilibrium states, is widely accepted as its mechanism. While the ETH has been numerically demonstrated to hold for various nonintegrable systems [11, 12], it cannot be applied to integrable systems due to the presence of local conserved quantities. Some experiments have also shown that integrable systems exhibit unusual relaxation dynamics [13]. Moreover, the existence of local conserved quantities is known to give rise to nontrivial inequalities in linear response [14, 15]. Thus, significant differences exist between the universal behavior observed in nonintegrable systems and the anomalous behavior in integrable systems. Consequently, in addressing the fundamental question of statistical mechanics—how the empirical laws of macroscopic systems emerge from quantum mechanics—it is crucial to determine whether the system is integrable or nonintegrable.
@@
While integrability has traditionally been a focus of rigorous research, nonintegrability has long remained beyond its scope. However, recent studies have developed methods to prove nonintegrability, or more specifically, the absence of local conserved quantities [16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27]. The initial work was limited to a spin-1/2 XYZ-h chain [16], but subsequent studies [17, 21, 22] have classified the nonintegrability of general translationally invariant spin-1/2 chains with nearest-neighbor symmetric interactions. This research has expanded to include spin-1/2 chains with next-nearest-neighbor interactions [18, 20, 27], spin-1 chains [23, 24], and higher-dimensional systems [25, 26]. On the other hand, the existing method is heuristic, relying on a form of “craftsmanship” to individually show the absence of $k$-local conserved quantities for each size $k$ of the operator’s support. Although all the proofs of nonintegrability in previous research have employed similar methods, the underlying unifying structure remains largely unexplored. This lack of a unified framework also poses a significant challenge when determining nonintegrability by using computers because the dimension of the space of $k$-local operators increases exponentially with $k$. Since other numerical tests for nonintegrability, such as the investigation of level spacing statistics [28, 29], are inherently affected by finite-size effects, determining whether the rigorous proof method described above can be formulated as an efficient algorithm is of fundamental importance.
@@
Efficient methods for determining integrability have long been of interest as well in the study of integrable systems [30, 31, 32, 33, 34]. It is conjectured that the existence of $3$-local conserved quantities is a necessary condition for integrability in one-dimensional systems with translational symmetry and nearest-neighbor interactions [34]. Conversely, previous proofs of nonintegrability of such systems [16, 17, 19, 21, 22, 23, 24] have supported the inverse  111 Strictly speaking, it is not literally inverse because there can exist systems that are neither integrable nor nonintegrable, but partially integrable, i.e., with some nontrivial local conserved quantities. Surprisingly, the existence of such partially integrable models is denied in the models examined in previous studies of nonintegrability. conjecture: the absence of $3$-local conserved quantities is a sufficient condition for nonintegrability [21]. However, these tests for integrability and nonintegrability remain empirical, and, to the best of our knowledge, no rigorous proof has yet been established. The main purpose of this paper is to present a rigorously provable test for integrability and nonintegrability based only on $3$-local quantities. This greatly simplifies the existing proof of nonintegrability and gives a partial solution to the above conjectures. Using our rigorous test, we construct an algorithm, which is independent of the system size, to guarantee nonintegrability of translationally invariant systems.
@@
This paper is organized as follows. In Sec. II, we fix the notation and state the main theorem. As an application of the theorem, we prove the absence of local conserved quantities in the XXZ chain with a non-uniform magnetic field. Section III is devoted to the proof of the theorems. Our discussion is based on some graphical language, which is also introduced in this section. Technical details of the proof are provided in Appendices. We also argue that our theorem and its proof can be adopted in the analysis of a related structure: spectrum generating algebras. In Sec. IV, we construct an algorithm of integrability and nonintegrability test for translationally invariant systems. We conclude the paper with a summary and outlook in Sec. V.
@@ II.1 Models
We consider a quantum spin chain on $N$ sites subject to the periodic boundary condition. The set of sites is denoted by $\Lambda_{N}=\{1,\dots,N\}$ and we identify $N+i$ with $i$. Each site has a state space $\mathcal{H}\cong\mathbb{C}^{d}$, which is a $d(<\infty)$-dimensional complex Hilbert space. Thus, the whole state space is $\mathcal{H}_{\mathrm{tot}}\coloneqq\mathcal{H}^{\otimes\Lambda_{N}}$. Note that the following argument holds for any finite $d$ and is not limited to spin-1/2 systems.
Next, we introduce $k$-local operators. The set of (bounded) linear operators and that of traceless operators on $\mathcal{H}$ are denoted by $\mathcal{B}$ and $\mathcal{B}_{0}$, respectively. For each operator $\hat{X}$ in $\mathcal{B}_{\mathrm{tot}}\coloneqq\mathcal{B}^{\otimes\Lambda_{N}}$, its support is defined as the smallest set of sites on which $\hat{X}$ acts nontrivially. The set of operators with a support $A(\subset\Lambda_{N})$ is defined as $\mathcal{B}_{0}^{A}\coloneqq\mathcal{B}_{0}^{\otimes A}$. Here we identify $\mathcal{B}^{\otimes A}$ with $\mathcal{B}^{\otimes A}\otimes I^{\otimes(\Lambda_{N}\setminus A)}$ where $I$ is the identity operator. The set of (strictly) $k$-local operators $\mathcal{B}^{(k)}$ is defined as a (direct) sum of $\mathcal{B}_{0}^{A}$ with $\operatorname{\mathrm{size}}(A)=k$, where
@@
is the size of a subset $A$. Note that $\mathcal{B}^{(0)}=\mathcal{B}_{0}^{\emptyset}=\mathbb{C}I^{\otimes\Lambda_{N}}$. We define the set of at most $k$-local quantities ~$\tilde{\mathcal{B}}^{(k)}$ as a (direct) sum of $\mathcal{B}^{(l)}$ for $0\leq l\leq k$. Since $\mathcal{B}=\mathcal{B}_{0}\oplus\mathbb{C}I$ holds, we have ~$\tilde{\mathcal{B}}^{(N)}=\mathcal{B}_{\mathrm{tot}}$. We define a length of an operator $\hat{X}$ as $\operatorname{\mathrm{len}}(\hat{X})\coloneqq\min\{k\in\{0,\dots,N\}\mid\hat{X}\in\tilde{\mathcal{B}}^{(k)}\}$, which represents how local $\hat{X}$ is. When we refer to the operator $\hat{X}$ simply as a $k$-local quantity, we mean $\operatorname{\mathrm{len}}(\hat{X})=k$. Such an operator can be decomposed into $l$-local terms for $0\leq l\leq k$: $\hat{X}=\sum_{l=0}^{k}\hat{X}^{(l)}$ with a nonvanishing $\hat{X}^{(k)}$, where $\hat{X}^{(l)}$ is an element of $\mathcal{B}^{(l)}$. If we emphasize that $\hat{X}^{(l)}$ is an element of $k$-local quantity $\hat{X}$, we write $\hat{X}_{[k]},\hat{X}_{[k]}^{(l)}$ instead of $\hat{X},\hat{X}^{(l)}$. For $l\geq 1$, each $\hat{X}^{(l)}$ is further expanded as $\hat{X}^{(l)}=\sum_{i=1}^{N}\hat{X}^{(l)}_{i}$, where
@@
Finally, we introduce a Hamiltonian of this system. We consider nearest-neighbor interactions and on-site potentials. This corresponds to a $2$-local Hamiltonian, which is represented as
@@
that is, $\hat{H}^{(2)}_{i}\in\mathcal{B}_{0}^{\{i,i+1\}}$ and $\hat{H}^{(1)}_{i}\in\mathcal{B}_{0}^{\{i\}}$. We are mainly interested in translationally invariant models, but we do not assume any relationship between these operators at this point. In addition, $\hat{H}$ can be non-hermitian unless otherwise stated.
@@ II.2 Main Result
$\hat{X}\in\mathcal{B}_{0}^{\{i+1\}},\ [\hat{H}^{(2)}_{i},I\otimes\hat{X}]=0$ $\Rightarrow\hat{X}=0$ (4)
$\mbox{and}\ \hat{Y}\in\mathcal{B}_{0}^{\{i\}},\ [\hat{H}^{(2)}_{i},\hat{Y}\otimes I]=0$ $\Rightarrow\hat{Y}=0$
for all $i\in\Lambda_{N}$, which ensures that this Hamiltonian is not “singular”. For example, let us consider an XYZ-type interaction: $\hat{H}^{(2)}_{i}=\sum_{\alpha=x,y,z}J^{\alpha}\hat{S}_{i}^{\alpha}\hat{S}_{i+1}^{\alpha}$ where $\hat{S}_{i}^{\alpha}$ represents the spin operator of spin $\mathcal{S}\in\mathbb{Z}_{>0}/2$. Then, for any $\mathcal{S}$, the assumption (4) is equivalent to the condition that at least two elements in $\{J^{x},J^{y},J^{z}\}$ are nonzero. Otherwise, this interaction becomes an Ising-type one or vanishes.
We are now in a position to state the main result of this paper: a rigorous test for integrability and nonintegrability. Without any further assumptions, we can show the following theorem rigorously.
@@
We assume a Hamiltonian $\hat{H}$ with nearest-neighbor interactions and on-site potentials (see Eq. (3)) satisfies the condition (4). If there is no $3$-local quantity $\hat{Q}$ such that $\operatorname{\mathrm{len}}([\hat{Q},\hat{H}])\leq 2$, then the system has no $k$-local conserved quantity for $23\leq k\leq N/$ 2. Equivalently, if there is a $k$-local conserved quantity for $23\leq k\leq N/$ 2, there also exists a $3$-local quantity $\hat{Q}$ such that $\operatorname{\mathrm{len}}([\hat{Q},\hat{H}])\leq 2$.
This theorem provides a sufficient condition for nonintegrability, or more precisely, the absence of $k$-local conserved quantities for $23\leq k\leq N/$ 2. Note that all models satisfying condition (4), among those whose nonintegrability has been proven [16, 19, 21, 22, 23, 24], satisfy this sufficient condition. As an equivalent consequence, Theorem 1 presents the existence of a 3-local quantity $\hat{Q}$ such that $\operatorname{\mathrm{len}}([\hat{Q},\hat{H}])\leq 2$ as a necessary condition for integrability. This yields a partial resolution of the conjecture in Ref. [34], which states that the existence of a 3-local conserved quantity $\hat{Q}$, that is, a 3-local quantity satisfying $\operatorname{\mathrm{len}}([\hat{Q},\hat{H}])=0$ is necessary for integrability. In summary, Thm. 1 gives a unified perspective and greatly simplifies existing proofs of nonintegrability, and provides a rigorous foundation for empirical tests of integrability. We note that Thm. 1 can be applied to models with finite-range interactions, spin ladders and models of higher-dimensional systems 222 In this case, the “locality” of an operator is defined in a very weak sense; it is determined by the size of its support for one direction. , because such a model is mathematically equivalent to some spin chain with a larger local dimension.
We can prove a more general theorem, which limits a possible form of conserved quantities. We introduce the following sets of operators:
@@
$\mathcal{B}_{<}^{(k)}$$\coloneqq\{\hat{X}\in\mathcal{B}_{\leq}^{(k)}\oplus\mathcal{B}^{(k-1)}\mid\operatorname{\mathrm{len}}([\hat{X},\hat{H}])\leq k-1\}.$ (6)
Note that, since $\operatorname{\mathrm{len}}([\hat{X},\hat{H}])\leq\operatorname{\mathrm{len}}(\hat{X})+1$ holds, $\hat{X}_{[k]}^{(k)}\in\mathcal{B}_{\leq}^{(k)}$ and $\hat{X}_{[k]}^{(k)}+\hat{X}_{[k]}^{(k-1)}\in\mathcal{B}_{<}^{(k)}$ are the easiest and second easiest conditions that must be satisfied for a $k$-local quantity $\hat{X}_{[k]}=\sum_{l=0}^{k}\hat{X}_{[k]}^{(l)}$ to be a conserved quantity. Under these definitions, the following theorem holds.

### v3

@@ Abstract
The integrability of a quantum many-body system, which is characterized by the presence or absence of local conserved quantities, drastically impacts the dynamics of isolated systems, including thermalization. Nevertheless, a rigorous and comprehensive method for determining integrability or nonintegrability has remained elusive. In this paper, we address this challenge by introducing rigorously provable tests for integrability and nonintegrability of quantum spin systems with finite-range interactions. Our results significantly simplify existing proofs of nonintegrability, such as those for the $S=1/2$ Heisenberg chain with nearest-and next-nearest-neighbor interactions, the $S=1$ bilinear-biquadratic chain and the $S=1/2$ XYZ model in two or higher dimensions. Moreover, our results also yield the first proof of nonintegrability for models such as the $S=1/2$ Heisenberg chain with a non-uniform magnetic field, the $S=1/2$ XYZ model on the triangular lattice, and the general spin XYZ model. This work also offers a partial resolution to the long-standing conjecture that integrability is governed by the existence of local conserved quantities with small support. Our framework ensures that the nonintegrability of one-dimensional spin systems with translational symmetry can be verified algorithmically, independently of system size.
@@ I Introduction
A central challenge in quantum many-body physics is to understand the equilibrium states and dynamics of strongly interacting systems. Rigorous methods sometimes offer deep insights into these problems. One such approach is to establish physically meaningful results for broad classes of systems, exemplified by the Hohenberg–Mermin–Wagner theorem [1, 2] and the Lieb–Schultz–Mattis theorem [3]. Another is the use of exactly solvable models. The Bethe ansatz method, originating from Bethe’s exact solution of the Heisenberg chain [4], enables the analytical calculation of quantities that are otherwise inaccessible in general systems, such as energy eigenvalues, eigenstates, and correlation functions. Furthermore, the algebraic Bethe ansatz method, or the quantum inverse scattering method [5], has revealed that behind such solvability lies integrability, i.e., the existence of an extensive number of local conserved quantities.
@@
However, when considering the dynamics of isolated quantum many-body systems, integrable systems can exhibit anomalous behavior. A notable example is their deviation from typical thermalization behavior [6]. Recent experiments have shown thermalization, i.e., the relaxation of local observables to thermal equilibrium in well-isolated quantum systems [7]. The eigenstate thermalization hypothesis (ETH) [8, 9, 10], which posits that energy eigenstates are locally indistinguishable from thermal equilibrium states, is widely accepted as the underlying mechanism. While the ETH has been numerically demonstrated to hold for various nonintegrable systems [11, 12], it cannot be applied to integrable systems due to the presence of local conserved quantities. Some experiments have also shown that integrable systems exhibit unusual relaxation dynamics [13]. Moreover, the existence of local conserved quantities is known to give rise to nontrivial inequalities in linear response [14, 15]. Thus, significant differences exist between the universal behavior observed in nonintegrable systems and the anomalous behavior in integrable systems. Consequently, in addressing the fundamental question of statistical mechanics—how the empirical laws of macroscopic systems emerge from quantum mechanics—it is crucial to determine whether the system is integrable or nonintegrable.
@@
While integrability has traditionally been a focus of rigorous research, nonintegrability has long remained outside its scope. However, recent studies have introduced methods for rigorously proving nonintegrability –or, more precisely, the absence of local conserved quantities [16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27]. The initial work focused on the spin-1/2 XYZ-h chain [16], but subsequent studies [17, 21, 22] classified the nonintegrability of general translationally invariant spin-1/2 chains with nearest-neighbor symmetric interactions. This research has expanded to include spin-1/2 chains with next-nearest-neighbor interactions [18, 20, 27], spin-1 chains [23, 24], and higher-dimensional systems [25, 26]. However, existing methods remain heuristic in nature, often relying on a form of “craftsmanship” to demonstrate the absence of $k$-local conserved quantities for each size $k$ of the operator’s support. Although all previous proofs of nonintegrability have employed similar methods, the underlying unifying structure remains largely unexplored. This lack of a unified framework also poses a significant challenge for computational approaches, as the dimension of the space of $k$-local operators grows exponentially with $k$. Since other numerical tests for nonintegrability, such as the investigation of level spacing statistics [28, 29], are inherently subject to finite-size effects, it is of fundamental importance to determine whether the rigorous proof method described above can be formulated as an efficient algorithm.
@@
Efficient methods for determining integrability have long been of interest as well in the study of integrable systems [30, 31, 32, 33, 34, 35]. It has been conjectured that the existence of $3$-local conserved quantities is a necessary condition for integrability in one-dimensional systems with translational symmetry and nearest-neighbor interactions [34, 35]. Conversely, previous proofs of nonintegrability in such systems [16, 17, 19, 21, 22, 23, 24] have supported the contrapositive 111 Strictly speaking, it is not literally contrapositive because there can exist systems that are neither integrable nor nonintegrable, but partially integrable, i.e., with some nontrivial local conserved quantities. Surprisingly, the existence of such partially integrable models is denied in the models examined in previous studies of nonintegrability. conjecture: the absence of $3$-local conserved quantities is a sufficient condition for nonintegrability [21]. However, these tests for integrability and nonintegrability remain empirical, and, to the best of our knowledge, no fully rigorous proof has yet been established. The main goal of this paper is to present a rigorously provable test for integrability and nonintegrability based solely on $3$-local quantities. Our approach unifies the conventional, handcrafted methods used in the study of nonintegrability and provides a partial resolution to the above conjectures. Using our rigorous test, we construct an algorithm–independent of system size–that guarantees the nonintegrability of translationally invariant quantum spin chains.
@@
This paper is organized as follows. In Sec. II, we fix the notation and state the main theorem. As an application of the theorem, we demonstrate the absence of local conserved quantities in several systems in Sec. III. Section IV is devoted to proving the theorems, where we also introduce a graphical language used in the discussion. Technical details of the proof are provided in Appendices. We also argue that our theorem and its proof can be applied to the analysis of a related structure: spectrum generating algebras. In Sec. V, we construct an algorithm to test integrability and nonintegrability in translationally invariant chains. Sec. VI presents an alternative test for systems where the method in Sec. II fails to apply. We conclude the paper with a summary and outlook in Sec. VII.
@@ II.1 Models
We consider a quantum spin chain on $N$ sites subject to the periodic boundary condition. The set of sites is denoted by $\Lambda_{N}=\{1,\dots,N\}$, where we identify $N+i$ with $i$. Each site has a local state space $\mathcal{H}\cong\mathbb{C}^{d}$, which is a $d$-dimensional complex Hilbert space with $d<\infty$. Accordingly, the total Hilbert space is $\mathcal{H}_{\mathrm{tot}}\coloneqq\mathcal{H}^{\otimes\Lambda_{N}}$. Note that the following argument apply to arbitrary finite $d$, and are not restricted to spin-1/2 systems.
Next, we introduce $k$-local operators. Let $\mathcal{B}$ denote the set of (bounded) linear operators on $\mathcal{H}$, and $\mathcal{B}_{0}$ the set of traceless operators on $\mathcal{H}$. For each operator $\hat{X}$ in $\mathcal{B}_{\mathrm{tot}}\coloneqq\mathcal{B}^{\otimes\Lambda_{N}}$, its support is defined as the smallest set of sites on which $\hat{X}$ acts nontrivially. The set of (traceless) operators supported on a subset $A(\subset\Lambda_{N})$ is defined as $\mathcal{B}_{0}^{A}\coloneqq\mathcal{B}_{0}^{\otimes A}$. Here we identify $\mathcal{B}_{0}^{\otimes A}$ with $\mathcal{B}_{0}^{\otimes A}\otimes I^{\otimes(\Lambda_{N}\setminus A)}$ where $I$ is the identity operator. The set of (strictly) $k$-local operators $\mathcal{B}^{(k)}$ is defined as a (direct) sum of $\mathcal{B}_{0}^{A}$ with $\operatorname{\mathrm{size}}(A)=k$, where
@@
is the size of a subset $A$. We set $\mathcal{B}^{(0)}=\mathcal{B}_{0}^{\emptyset}=\mathbb{C}I^{\otimes\Lambda_{N}}$. We define the set of at most $k$-local quantities ~$\tilde{\mathcal{B}}^{(k)}$ as a (direct) sum of $\mathcal{B}^{(l)}$ for $0\leq l\leq k$. Since $\mathcal{B}=\mathcal{B}_{0}\oplus\mathbb{C}I$ holds, we have ~$\tilde{\mathcal{B}}^{(N)}=\mathcal{B}_{\mathrm{tot}}$. We define the length of an operator $\hat{X}$ as $\operatorname{\mathrm{len}}(\hat{X})\coloneqq\min\{k\in\{0,\dots,N\}\mid\hat{X}\in\tilde{\mathcal{B}}^{(k)}\}$, which quantifies the degree of locality of $\hat{X}$. When we refer to the operator $\hat{X}$ simply as a $k$-local quantity, we mean that $\operatorname{\mathrm{len}}(\hat{X})=k$. Such an operator can be decomposed into $l$-local terms for $0\leq l\leq k$: $\hat{X}=\sum_{l=0}^{k}\hat{X}^{(l)}$ with a nonvanishing $\hat{X}^{(k)}$, where $\hat{X}^{(l)}$ is an element of $\mathcal{B}^{(l)}$. If we emphasize that $\hat{X}^{(l)}$ is a component of $k$-local quantity $\hat{X}$, we write $\hat{X}_{[k]},\hat{X}_{[k]}^{(l)}$ instead of $\hat{X},\hat{X}^{(l)}$. For $l\geq 1$, each $\hat{X}^{(l)}$ is further expanded as $\hat{X}^{(l)}=\sum_{i=1}^{N}\hat{X}^{(l)}_{i}$, where
@@
Finally, we introduce the Hamiltonian of the system, which consists of nearest-neighbor interactions and on-site potentials. This corresponds to a $2$-local Hamiltonian, which is represented as
@@
that is, $\hat{H}^{(2)}_{i}\in\mathcal{B}_{0}^{\{i,i+1\}}$ and $\hat{H}^{(1)}_{i}\in\mathcal{B}_{0}^{\{i\}}$. We are mainly interested in translationally invariant systems, but we do not assume any relationship between these operators at this point. In addition, $\hat{H}$ can be non-hermitian.
@@ II.2 Main Result
$\hat{X}\in\mathcal{B}_{0}^{\{i+1\}},\ [I\otimes\hat{X},\hat{H}^{(2)}_{i}]=0$ $\Rightarrow\hat{X}=0$ (4)
$\mbox{and}\ \hat{Y}\in\mathcal{B}_{0}^{\{i\}},\ [\hat{Y}\otimes I,\hat{H}^{(2)}_{i}]=0$ $\Rightarrow\hat{Y}=0$
for all $i\in\Lambda_{N}$. We refer to this condition as injectivity, which ensures that this Hamiltonian is not “singular”. For example, let us consider an XYZ-type interaction: $\hat{H}^{(2)}_{i}=\sum_{\alpha=x,y,z}J_{\alpha}\hat{S}_{i}^{\alpha}\hat{S}_{i+1}^{\alpha}$ where $\hat{S}_{i}^{\alpha}$ represents the spin operator of spin §$\S\in\mathbb{Z}_{>0}/2$. Then, for any §$\S$, the injectivity (4) is equivalent to requiring that at least two elements in $\{J_{x},J_{y},J_{z}\}$ are nonzero. Otherwise, this interaction becomes an Ising-type one or vanishes.
We are now in a position to state the main result of this paper: a rigorous test for integrability and nonintegrability. We first state a result that holds under an additional assumption that is satisfied in many systems of interest. We introduce the following set of operators:
$\mathcal{B}_{\leq}^{(2)}$$\coloneqq\{\hat{X}\in\mathcal{B}^{(2)}\mid\operatorname{\mathrm{len}}([\hat{X},\hat{H}])\leq 2\}$
=$=\{\hat{X}\in\mathcal{B}^{(2)}\mid\operatorname{\mathrm{len}}([\hat{X},\hat{H}^{(2)}])\leq 2\}.$ (5)
It is evident that $\mathcal{B}_{\leq}^{(2)}$ contains $\hat{H}^{(2)}$. Conversely, for generic spin chains with sufficiently complex interactions, it is expected that $\mathcal{B}_{\leq}^{(2)}$ is limited to this, i.e.,
$\mathcal{B}_{\leq}^{(2)}=\mathbb{C}\hat{H}^{(2)}$ (6)
holds. For example, the spin-$S$ XYZ chains satisfy Eq. (6) for arbitrary $S\in\mathbb{Z}_{>0}/2$; see the last paragraph of Sec. V and Appx. C. Under these assumptions, we can show the following theorem rigorously.
@@
Consider a Hamiltonian $\hat{H}$ with nearest-neighbor interactions and on-site potentials (see Eq. (3)) satisfying the assumptions (4) and (6). If there is no $3$-local quantity $\hat{Q}$ such that $\operatorname{\mathrm{len}}([\hat{Q},\hat{H}])\leq 2$, then the system has no $k$-local conserved quantity for any $23\leq k\leq N/$ 2. Equivalently, if there is a $k$-local conserved quantity for some $23\leq k\leq N/$ 2, there also exists a $3$-local quantity $\hat{Q}$ such that $\operatorname{\mathrm{len}}([\hat{Q},\hat{H}])\leq 2$.
This theorem provides a sufficient condition for nonintegrability, or more precisely, the absence of $k$-local conserved quantities for $23\leq k\leq N/$ 2. Note that all models satisfying condition (4), among those whose nonintegrability has been proven [16, 19, 21, 22, 23, 24], satisfy this sufficient condition. Equivalently, Thm. 1 implies that the existence of a 3-local quantity $\hat{Q}$ with $\operatorname{\mathrm{len}}([\hat{Q},\hat{H}])\leq 2$ is a necessary condition for integrability. This yields a partial resolution of the conjecture in Ref. [34], which states that the existence of a 3-local conserved quantity $\hat{Q}$, that is, a 3-local quantity satisfying $\operatorname{\mathrm{len}}([\hat{Q},\hat{H}])=0$ is necessary for integrability. In summary, Thm. 1 gives a unified perspective and greatly simplifies existing proofs of nonintegrability, and provides a rigorous foundation for empirical tests of integrability.
Thm. 1 can be applied to models with finite-range interactions, because such a model is mathematically equivalent to some spin chain with a larger local dimension. Note that there is an extension of the above conjecture to medium-range interactions rather than relying on such a transformation to nearest-neighbor interactions [35]. We can also regard spin ladders and systems on higher-dimensional lattices as one-dimensional systems in the same way 222 In this case, the “locality” of an operator is defined in a very weak sense; it is determined by the size of its support for some direction. ; see, e.g., Subsec. III.2. While Eq. (6) is expected to be satisfied if the interaction is sufficiently complex, it does not hold for models with nearest-neighbor interactions on the hypercubic lattice or the honeycomb lattice. Instead, we will explain a method to simplify the proof of nonintegrability in such systems in Sec. VI.
Next, we consider general systems with the injectivity, which may not satisfy the assumption (6). We can prove some results even in this case, which limit a possible form of conserved quantities. As generalizations of $\mathcal{B}_{\leq}^{(2)}$, we introduce the following sets of operators:
@@
$\mathcal{B}_{<}^{(k)}\coloneqq\{\hat{X}\in\mathcal{B}^{(k)}_{\leq}\oplus\mathcal{B}^{(k-1)}\mid\operatorname{\mathrm{len}}([\hat{X},\hat{H}])\leq k-1\}.$ (8)
If it is necessary to indicate the dependence on $\hat{H}$, we denote $\mathcal{B}_{\leq}^{(k)}$ and $\mathcal{B}_{<}^{(k)}$ by $\mathcal{B}_{\leq}^{(k)}(\hat{H})$ and $\mathcal{B}_{<}^{(k)}(\hat{H})$, respectively. Note that, since $\operatorname{\mathrm{len}}([\hat{X},\hat{H}])\leq\operatorname{\mathrm{len}}(\hat{X})+1$ holds, $\hat{X}_{[k]}^{(k)}\in\mathcal{B}_{\leq}^{(k)}$ and $\hat{X}_{[k]}^{(k)}+\hat{X}_{[k]}^{(k-1)}\in\mathcal{B}_{<}^{(k)}$ are the easiest and second easiest conditions that must be satisfied for a $k$-local quantity $\hat{X}_{[k]}=\sum_{l=0}^{k}\hat{X}_{[k]}^{(l)}$ to be a conserved quantity. For $\mathcal{B}_{\leq}^{(k)}$’s, we have the following theorem.